steps to solve x^4 + 2x^2-15=0
Found 2 solutions by lwsshak3, stanbon:
Answer by lwsshak3(11628)

steps to solve x^4 + 2x^2-15=0
factor
(x^2+5)(x^2-3)=0
x^2+5=0
x^2=-5
x=±√-5
x=±(√5) i
or
x^2-3=0
x^2=3
x=±√3
two real roots: ±√3
two imaginary or non-real roots: ±(√5) i

Answer by stanbon(75887)

steps to solve x^4 + 2x^2-15=0
-----
Factor:
(x^2+5)(x^2-3) = 0
---
x^2 = -5 or x^2 = 3
---
Real-Number solutions:
x = sqrt(3) or x = -sqrt(3)
============================
Cheers,
Stan H.
=====================